### v2

@@
Abstract. We develop an NLP-based procedure for detecting systematic nonmeritorious consumer complaints, simply called systematic anomalies, among complaint narratives. While classification algorithms are used to detect pronounced anomalies, in the case of smaller and frequent systematic anomalies, the algorithms may falter due to a variety of reasons, including technical ones as well as natural limitations of human analysts. Therefore, as the next step after classification, we convert the complaint narratives into quantitative data, which are then analyzed using an algorithm for detecting systematic anomalies. We illustrate the entire procedure using complaint narratives from the Consumer Complaint Database of the Consumer Financial Protection Bureau.
@@ 1 Introduction
The narratives of consumer complaints contain valuable information about the sales and underwriting activities in the insurance industry. Responding to consumer complaints is one of the primary ways for insurance departments (NAIC, 2023) to regulate market practices (Klein, 1995). In 2021, state insurance departments received 259,345 official complaints, among which 1,474 market conduct examinations were triggered and completed (NAIC, 2022). By analyzing consumer complaints data, regulators can identify market mis-behaviour, such as false sales illustrations or failure to pay legitimate claims on a timely basis, and initiate market conduct examinations to determine if disciplinary procedures are needed.
@@
Certainly, datasets such as the CFPB consumer complaints also serve the purpose of regulation, as using complaints for regulatory purposes is significantly more cost-effective than resolving them individually (Littwin, 2014). Analysts have attempted to determine whether complaints have merits and possibly constitute violations of state laws or regulations (Klein, 1995). In such cases, companies may voluntarily provide reliefs along with responses to the complaints. These reliefs can be seen as risk premiums to mitigate regulatory or litigation risks in the cases described by complaint narratives. Hence, reliefs, particularly the tangible ones, are critical indicators of potential violations of state laws or regulations. Among the complaints in the CFPB database, companies have granted consumers tangible reliefs 21.9% of the time. In the remaining cases, the focus has been on providing consumers with explanations (Littwin, 2014). Nevertheless, whether intentionally or not, companies are reluctant to grant reliefs in some cases, which we call anomalies in this paper.
@@
For those datasets that include the narratives of complaints, the Natural Language Processing (NLP) based analyses have become popular. For example, Karami and Pendergraft (2018) propose a computational approach for analyzing complaints from an insurance company with more than 16 million policies and 24 million insured vehicles. Using a similar method, Liao et al. (2020) develop a process for an insurance company of personal lines to improve its customer services by analyzing customer calls. Furthermore, based on the CFPB dataset, Osman and Sabit (2022) investigate the impact of a major banking scandal using the Valence Aware Dictionary for sEntiment Reasoning (VADER) sentiment intensities (Hutto and Gilbert, 2014). Pakhchanyan et al. (2022) provide an overview of how machine learning helps to categorize textual descriptions of operational loss events into Basel II event types by applying PYTHON implementations of support vector machine (SVM) and multinomial naive Bayes algorithms to pre-categorized Öffentliche Schadenfälle OpRisk (ÖffSchOR) data. Wang et al. (2022) introduce a text mining method to analyze changes in operational risk in an innovative way based on a massive data set of textual risk disclosures of financial institutions, which aggregate risk perceptions of senior managers of the financial industry. Ji et al. (2023) develop a risk-based knowledge graph to integrate text mining and analytic hierarchy process risk assessment with knowledge graphs for operational risk analysis.
@@
In the current paper, we focus on developing a systematic-anomaly detection procedure for assessing complaints for which companies are reluctant to provide reliefs. In our study, the VADER sentiment intensities play a crucial role in the text classifier training stage as well as in the performance evaluation stage. In the text classifier stage, the VADER sentiment intensities are combined with the Term Frequency-Inverse Document Frequency (TF-IDF) for the purpose of feature extraction. In the performance evaluation stage, the VADER sentiment intensities are combined with the discounted dollar amounts mentioned in the complaint narratives as well as with the word counts of the cleaned narratives. This, in turn, facilitates the use of anomaly detection indices, which test if there are persistent anomalies in the dataset of complaints. To illustrate the proposed methodology, we use the CFPB dataset due to its transparency and accessibility.
The rest of this paper is organized as follows. In Section 2, we describe a subset of the CFPB consumer complaints data. In Section 3, we focus on anomaly detection among complaint narratives using classification algorithms. In Section 4, we convert the narratives into quantitive data and establish a basis for input-output systems. In Section 5, we define the input-output systems arising from two featurizations. In Section 6, we introduce anomaly detection indices based on the input-output systems described in the previous sections, and then evaluate the ability of these indices to detect anomalies. Section 7 concludes the paper.
@@ 2 Data
We analyze a subset of complaint narratives from the Consumer Complaint Database of the Consumer Financial Protection Bureau (CFPB, 2023). Namely, we extract the records of those complaints whose narratives were submitted between December 1, 2011, and June 29, 2023. We focus on the complaints related to credit cards and prepaid cards of a major bank. Of all the selected complaint records, only those referring to unique dollar amounts are considered. Furthermore, we exclude the complaints referring to dollar amounts exceeding \$10,000 because those complaints are more likely to be associated with business accounts rather than individual accounts. The resulting data consist of 2,905 records of complaints related to credit card and prepaid card services of the bank. For simplicity, we select the complaints with only single positive dollar amounts appearing in the narratives. As a consequence, we obtain 2,894 complaints for our study.
The complaints closed without any (monetary or non-monetary) reliefs are viewed as nonmeritorious (Hayes et al., 2021), which we call anomalies throughout the paper. We note that among the 2,894 complaints, 1,051 were meritorious, that is, they received some reliefs. As there are dollar amounts in our selected complaint narratives submitted at different times, we discount the amounts using the annual CPI (CPI, 2023) to the base date of January 1, 2015. Table 2.1
Date Narrative Adjusted Merit

### v3

@@
Abstract. A Natural Language Processing (NLP) based procedure is developed for detecting systematic non-meritorious consumer complaints, simply called systematic anomalies, among complaint narratives. While classification algorithms are used to identify meritorious complaints, in the case of smaller and frequent systematic patterns of non-meritorious complaints, those algorithms may falter due to a variety of reasons, including technical ones as well as natural limitations of human analysts. Therefore, at the next stage after classification, the complaint narratives are converted into quantitative data, which are then analyzed using indices for detecting systematic anomalies. An illustration of the entire procedure is provided using complaint narratives from the Consumer Complaint Database of the Consumer Financial Protection Bureau. The results suggest that the Support Vector Machine (SVM) outperforms other selected classifiers. Although classification results with the VADER intensity pertinent to the featurization step have lower accuracy, they contain fewer non-meritorious complaints than those without the VADER intensity.
@@ 1 Introduction
The narratives of consumer complaints contain valuable information about the sales and underwriting activities in the insurance industry. Responding to consumer complaints is one of the primary ways for insurance departments (NAIC, 2023) to regulate market practices (Klein, 1995). In 2021, state insurance departments received 259 345 official complaints, among which 1 474 market conduct examinations were triggered and completed (NAIC, 2022). By analyzing consumer complaint data, regulators can identify market mis-behaviour, such as false sales illustrations or failure to pay legitimate claims on a timely basis, and initiate market conduct examinations to determine if disciplinary procedures are needed.
@@
Certainly, datasets such as the CFPB consumer complaints also serve the purpose of regulation, as using complaints for regulatory purposes is significantly more cost-effective than resolving them individually (Littwin, 2014). Analysts have attempted to determine whether complaints have merits and possibly constitute violations of state laws or regulations (Klein, 1995). In such cases, companies may voluntarily provide reliefs along with responses to the complaints. These reliefs can be seen as risk premiums to mitigate regulatory or litigation risks in the cases described by complaint narratives. Hence, reliefs, particularly the tangible ones, are critical indicators of potential violations of state laws or regulations. Among the complaints in the CFPB database, companies have granted consumers tangible reliefs 21.9% of the time. In the remaining cases, the focus has been on providing consumers with explanations (Littwin, 2014). Nevertheless, whether intentionally or not, companies are reluctant to grant reliefs in some cases, which are viewed in the present paper as anomalies of the dispute-resolving system.
1.1 Motivation
Reliefs to consumer complaints should be granted with prudence. From the economic perspective, reliefs to consumer complaints are costly from the time, effort, and monetary perspectives, and hence impact the profitability of businesses. Moreover, high percentages of reliefs granted to complaints suggest poor levels of compliance with regulations or operational risk management, and hence influence the reputation of financial institutions by attracting fewer customers although more media attention and regulation. For this reason, financial institutions need to implement operational risk management measures to control the percentage of complaints that deserve reliefs. Indeed, taking risks brings opportunities, and so completely avoiding risks to minimize the aforementioned percentage may not be attractive from the business perspective. Yet, the percentage cannot be too high due to reasons such as arising reputational and regulatory issues.
With the help of records of monetary and non-monetary reliefs granted to complaints provided by the publicly available CFPB dataset, machine learning based classifiers can be trained to assist in determining whether reliefs should be granted. Among the complaints submitted to the CFPB, those that are closed with monetary or non-monetary reliefs are defined as “meritorious” (Hayes et al., 2021) throughout the present paper .
There are several reasons to suggest that more effort is needed for this classification task. First, unlike typical classification tasks of text inputs, classifiers cannot be trained by using comments along with polarities (e.g., tweets) or the number of stars (e.g., online shopping experiences) as responses. Second, complaints are submitted by consumers whereas decisions to grant, or not to grant, reliefs are made by specialists of financial institutions. As a result, the potential features extracted from the narratives are often less correlated with the responses, thus leading to relatively low accuracy if compared with the polarity classification of tweets or online shopping experiences. Third, extra steps are needed to check if more complaints can actually be filtered out from the pool of meritorious complaints.
1.2 Related work
@@
For those datasets that include the narratives of complaints, the Natural Language Processing (NLP) based analyses have become popular. For example, Karami and Pendergraft (2018) propose a computational approach for analyzing complaints from GEICO, an insurance company with more than 16 million policies and 24 million insured vehicles. However, only the Latent Dirichlet Allocation (LDA) based topic analysis is performed such that 30 categories of topics are identified without any direct relationship to decisions of pay-offs or reliefs. Using a similar method, Liao et al. (2020) consider the LDA-based topic analysis for an insurance company of personal lines to improve its customer services by analyzing customer calls. As the LDA-based topic analysis is an unsupervised learning technique, this study does not use decisions related to payoffs or reliefs as classification labels, although it involves other NLP techniques such as text featurization and sentiment analysis. Furthermore, based on the CFPB dataset, Osman and Sabit (2022) investigate the impact of a major banking scandal using the Valence Aware Dictionary for sEntiment Reasoning (VADER) (Hutto and Gilbert, 2014). However, this study directly quantifies text inputs using VADER without a further exploration of text mining or classification problems. Pakhchanyan et al. (2022) provide an overview of how machine learning helps to categorize textual descriptions of operational loss events into the seven Basel II event types of operational risk by applying PYTHON implementations of the SVM and multinomial naive Bayes algorithms to pre-categorized Öffentliche Schadenfälle OpRisk (ÖffSchOR) data.
In the present study, the inputs are the textual descriptions of operational loss events (e.g., Di Vincenzo et al., 2023), which have higher quality due to professional narratives compared with consumer complaints that contain more typos but less informative details. Wang et al. (2022) introduce a text mining method to analyze changes in operational risk in an innovative way based on a massive data set of textual risk disclosures of financial institutions, which aggregate risk perceptions of senior managers of the financial industry. However, this study employs textual data analytic techniques to identify significant operational risk drivers instead of performing machine learning tasks such as classification. Ji et al. (2023) develop a risk-based knowledge graph to integrate text mining and analytic hierarchy process risk assessment with knowledge graphs for operational risk analysis. While text mining techniques are integrated in this study, quantifying operational risk rather than assisting decision-making is still the ultimate goal (e.g., Wang et al., 2022; Osman and Sabit, 2022). Di Vincenzo et al. (2024) provide a tweet data analysis for detecting emerging operational risks.
1.3 Research gaps
Due to poor quality of narratives as textual data inputs as well as weak correlation between the inputs (from customers) and outputs (from financial institutions), the existing text classification methods have shortcomings, particularly from the perspective of limiting the percentage of meritorious complaints. These considerations have motivated us to work on improving classification procedures from several perspectives, including the featurization of textual data inputs, and to also propose additional analyses using an anomaly detection method to assess if the resulting meritorious classes actually contain systematic non-meritorious complaints.
1.4 The scope
@@
In the present paper, a systematic-anomaly detection procedure is proposed to assess complaints for which companies are reluctant to provide reliefs. In the current study, the VADER sentiment intensities play an important role in the text classifier training stage as well as in the performance evaluation stage. In the text classifier stage, the VADER sentiment intensities are combined with the Term Frequency-Inverse Document Frequency (TF-IDF) for the purpose of feature extraction. In the performance evaluation stage, the VADER sentiment intensities are combined with the discounted dollar amounts mentioned in the complaint narratives as well as with the word counts of the cleaned narratives. This, in turn, facilitates the use of anomaly detection indices, which test if there are persistent anomalies in the dataset of complaints. To illustrate the proposed methodology, we use the CFPB dataset, due to its transparency and accessibility.
The present study is comprised of several stages. At the classification stage, the labels are taken directly from the CFPB complaint dataset, with the complaints that were resolved by either monetary or non-monetary reliefs labeled as “meritorious,” while complaints with only responses labeled as “non-meritorious” (Hayes et al., 2021). The selected complaints are then restricted to those consumer complaints that involve specific dollar amounts related to personal credit or debit card accounts against a selected major financial institution, specifically, the Bank of America (BOA). The features of the classification task are extracted from the complaint narratives pre-processed by NLP and transformed to numerical values using the widely utilized TF-IDF method and our modified version of the TD-IDF using the VADER sentiment density. The classifiers considered in this study include the Logistic Regression (LR), Support Vector Machine (SVM), Gradient Boosting (GB), Multilayer Perceptron (MLP), and Random Forest (RF), which are widely used in text classification studies (e.g., Hsu, 2020). At the performance evaluation stage, the resulting meritorious class is further assessed by the index for systematic-anomaly detection (Gribkova and Zitikis, 2018, 2019a, 2019b, 2020).
1.5 Contributions
The present study has several important implications on how financial institutions can better utilize massive volumes of customer complaints in order to train classifiers that support decision making when resolving disputes. Among the achievements are:
  1. (1)
    Exploring the possibility of improving textual data classification by incorporating the VADER sentiment density.
  2. (2)
    Introducing an anomaly-detection method that controls the percentage of reliefs granted to complaints.
  3. (3)
    Evaluating classifiers by examining the extent at which the resulting meritorious class includes actually non-meritorious complaints in a systematic pattern.
  4. (4)
    Demonstrating the usefulness of the proposed procedure in identification of meritorious customer complaints in the CFPB dataset.
The rest of the paper is organized as follows. In Section 2, a subset of the CFPB consumer complaint data is chosen for illustration of the proposed procedure. In Section 3, selected classification algorithms are employed for anomaly detection among complaint narratives. In Section 4, the narratives are converted into quantitative data, and a basis for an input-output system is established. In Section 5, two input-output systems arising from two featurization methods are introduced. In Section 6, anomaly detection indices based on the two input-output systems are introduced, and their ability to detect anomalies is evaluated. Section 7 concludes the paper. Two appendices illustrating the steps of corpus cleaning (e.g., Di Vincenzo et al., 2023) as well as the behaviour of the $B$-index with respect to increasing sample sizes are in the online supplement to this paper.
@@ 2 Data
A subset of complaint narratives from the Consumer Complaint Database of the Consumer Financial Protection Bureau (CFPB, 2023) is selected for illustrative purposes. Extracted are the records of those complaints whose narratives
  1. (1)
    are related to credit cards and prepaid cards of a major bank,
  2. (2)
    were submitted between December 1, 2011, and June 29, 2023.
Of all the selected complaint records, only those referring to unique dollar amounts are considered. Furthermore, the complaints referring to dollar amounts exceeding \$10 000 are excluded because those complaints are more likely to be associated with business accounts rather than individual accounts. Hence, the resulting data consist of 2 905 records of complaints related to credit card and prepaid card services of the bank. For simplicity, complaints with only single positive dollar amounts appearing in the narratives are selected. As a consequence, 2 849 complaints are ultimately selected for the current study.
The complaints closed without any monetary or non-monetary reliefs are viewed as non-meritorious (Hayes et al., 2021) and called anomalies throughout the present paper. Among the selected 2 849 complaints, 1 051 are meritorious, that is, they were resolved with some reliefs. As there are dollar amounts submitted at different times, we discount the amounts using the annual Consumer Price Index (CPI), which is defined (CPI, 2023)0
[…]
to the base date of January 1, 2015. Table 2.1
Date Raw complaint narratives Adjusted Merit